# On Scaling LT-Coded Blockchains in Heterogeneous Networks and their Vulnerabilities to DoS Threats

### v2

@@ 1 Introduction and Background
[…]
@@ 1.1 Our Contributions
Overall, our work is the first of its kind in designing optimal decoders for heterogeneous network constraints and also analysing the threats caused by non-oblivious adversaries in LT coded blockchain architecture.
@@ 1.2.1 Blockchain Scalability using Error Correction Codes
Figure 1: Number of works that implement RS, Fountain, LDPC and LCC codes to address scalability issues in blockchains (source: [14]).
@@
Furthermore, as another interesting direction of research, several coded architectures have been proposed to address the scalability issue in blockchains. In this context, authors of [17] present a secure fountain (SeF) architecture for blockchains by using LT codes to provide storage benefits. Ultimately, [17] aims to achieve a near-optimal trade-off between bootstrap overhead and storage savings. Authors of [19] have also worked on similar lines as [17], and suggested raptor codes to be used in place of LT codes to achieve an almost constant decoding complexity, however, with a higher bootstrap overhead. Subsequently, [32] pointed out the problem of high bandwidth requirement for decoding a particular requested block, faced by [17, 19]. To this end, [32] presented two different architectures namely (i) Secure and Private Fountain codes, (ii) Secure and Private Randomly Sampled codes to obtain storage savings, low repair bandwidth per epoch, low communication cost for decoding a requested block, and simultaneously provide privacy of the archival data. However, unlike [17, 19], the architectures in [32] incur higher costs for decoding the complete chain. Another work [33], presents a secure-repair-block protocol based on regenerating codes for sharded blockchains to reduce both storage and bootstrap costs, which makes it easier for a new miner to enter the system. A different study [34] introduces homomorphic hashes in coded blockchains, which allows instantaneous detection of erroneous fragments, thereby avoiding decoding with wrong data.
@@ 1.2.2 Data Availability Attacks in Blockchains
Some preliminary results of this work are available in [44], which captures a part of our contributions on the vulnerabilities of LT coded blockchains to optimized DoS threats. In particular, this work builds upon the work in [44] by providing a comprehensive analysis of LT decoders’ performance metrics and exploring the trade-off between computational complexity and bootstrap overhead, which [44] does not address. While [44] briefly discusses hybrid decoders, this work optimizes them for blockchain decoding at minimal cost under heterogeneous network constraints. In vulnerability analysis, we introduce more advanced and powerful attack strategies and efficient implementation techniques. Specifically, we propose a more effective score-based attack strategy on the BP decoder and devise a computationally efficient algorithm for attacking the OFG decoder, improving upon the attacks presented in [44].
@@ 2 LT Coded Blockchain Architecture
To accommodate the rapid growth in the size of a full node, coded blockchain architecture distributes the contents of a full node as coded fragments across multiple smaller-sized storage devices. Fig. 2 shows a visual illustration of coded blockchain architecture. In this architecture, the original blockchain stored on a full node is divided into several epochs, wherein an epoch is defined as a collection of $k$ blocks of the blockchain, for some $k\in\mathbb{Z}_{+}$. For every epoch, a sufficient number of coded blocks referred to as droplets is generated, and these droplets are stored across storage constrained nodes referred to as droplet nodes, as seen in Fig. 2. Henceforth, we assume that a droplet node will store one droplet per epoch, thus storing as many droplets as the number of epochs, as shown clearly in Fig. 2. This way, the storage size of a full node is slashed by distributing it across smaller-sized storage devices with roughly $\frac{1}{k}$ storage capacity. In the next section, we explain how a blockchain is encoded to obtain the droplet nodes using an LT coded architecture [17].
@@ 2.1 Slashing Storage Costs through LT Encoding
When generating the droplets of a full node, LT encoding is done as follows. First, a number $d\in[k]$ is chosen randomly from a suitable probability distribution. Subsequently, $d$ distinct input blocks are uniformly chosen from the $k$ blocks of an epoch, and a bit-wise XOR of the chosen blocks, denoted by $C$, is obtained. Henceforth, this collection of $d$ input blocks of an epoch that are XORed to obtain $C$ is referred to as the neighbours of the droplet $C$. Along with $C$, a binary vector $\mathbf{v}$ of length $k$ is also generated, wherein the $b$-th entry of $\mathbf{v}$ is $1$ if the $b$-th input block is among the $d$ chosen blocks, else the $b$-th entry is 00. This combination of $C$ and $\mathbf{v}$ constitute a droplet for an epoch. An illustration of this procedure is given in Fig. 3, for the case when $k=6$.
@@
Furthermore, we assume that all the droplets of a droplet node that correspond to different epochs, will have the same degree and also the same set of neighbours. In the LT code terminology, the number $d$, which is referred to as the degree, is picked from the popular Robust Soliton Distribution (RSD) [18, Definition 11], denoted by $\Omega_{RS}(\cdot)$ which is defined for a given set of parameters $10<\delta<$ 1 and $c>0$.
$\Omega_{RS}(d)=\frac{\rho(d)+\tau(d)}{\beta}\quad\forall d=1,\ldots,k,$
@@
$\rho(d)=\begin{cases}1/k&\text{for }d=1\\ \frac{1}{d(d-1)}&\text{for }d=2,\ldots,k,\\ \end{cases}$
@@
$\tau(d)=\begin{cases}\frac{R}{dk}&\text{for }d=1,\ldots,\frac{k}{R}-1\\ \frac{R}{k}\cdot\ln{\left(\frac{R}{\delta}\right)}&\text{for }d=\frac{k}{R}\\ 0&\text{for }d=\frac{k}{R}+1,\ldots,k,\end{cases}$
@@ 2.2.3 Hybrid Decoder
[…]
@@ 3.3 Performance Analysis
[…]
Overhead_BP The average bootstrap overhead of a bucket node employing the BP decoder, to successfully decode an epoch with a probability greater than $1-\delta$ is given by $K_{BP}=k+c\sqrt{k}\ln^{2}{\left(\frac{k}{\delta}\right)}$, where $c$ is a parameter of RSD.
@@
[…]
Computation_BP The average computational complexity incurred by the BP decoder for decoding an epoch is given by $C_{BP}=k\cdot\left(\ln\left(\frac{c\cdot k^{\frac{3}{2}}\cdot\ln(k/\delta)}{\delta}\right)+1.577\right)$.
@@
The computational complexity for decoding an epoch can be determined by the number of edges in the bipartite graph $\mathcal{T}$, generated at the start of the BP decoder [17, Appendix B]. This is because, each XOR operation between two droplets corresponds to removal of an edge in the bipartite graph. Subsequently, all the edges of the bipartite graph will be removed at the end of BP decoding process, thereby resulting in as many number of XOR operations. It is important to note that the number of edges in the bipartite graph is random in nature, by the virtue of the random selection of the droplet nodes. Therefore, we calculate its mean value to determine the average computational complexity. In this regard, the average number of edges can be approximated as the product of average bootstrap overhead and the average degree of a droplet. In particular, if $D$ is the average degree of a droplet, it is straightforward to show from [18, Theorem 13] that
$D=\frac{1}{\beta}\cdot\left(\ln\left(\frac{c\cdot k^{\frac{3}{2}}\cdot\ln(k/\delta)}{\delta}\right)+1.577\right),$
where $c$, $\delta$ and $\beta$ are parameters of RSD. Now, we can calculate the average computational complexity as $C_{BP}=K_{BP}D$. ∎
The above two results present closed-form expressions on the average bootstrap overhead and average computational complexity of the BP decoder.
Now, to derive the bootstrap overhead of OFG decoder, we must investigate the relationship between its probability of decoding failure and the number of downloaded droplets per epoch. To this end, authors of [46] have established an upper bound on the decoding failure probability, denoted by ${P_{F}}$, for an LT code whose degree follow any arbitrary degree distribution. As a result, for binary LT codes whose degrees follow RSD, the upper bound in [46, Theorem 1] can be modified as follows:
@@
$\cdot\left[\frac{1}{2}\sum_{d}\Omega_{RS}(d)\frac{\sum_{l=0}^{d}{w\choose l}{k-w\choose d-l}\left[1-(-1)^{1-l}\right]}{{k\choose d}}\right]^{k(1+\epsilon)},$
[…]
@@
[…]
Overhead_OFG Let $\epsilon_{min}$ = $\frac{m}{k}$ where $m\in\mathbb{N}$ such that $P_{F}|_{\epsilon=\epsilon_{min}}\leq\delta$, then the bootstrap overhead of OFG decoder is upper bounded by $K_{OFG}=k(1+\epsilon_{min})$.
@@
[…]
Computation_OFG Computational complexity of OFG decoder, denoted by $C_{OFG}$ is upper bounded by $k^{2}$.
@@
Computational complexity of the OFG decoder is evaluated as the sum of number of XOR operations performed by the OFG algorithm and back substitution, which together completes the decoding process. In the OFG algorithm, the maximum number of XOR operations associated with the $i^{th}$ droplet, whose binary coefficient entering the partially filled $k\times k$ matrix, is $i$. Hence, to fill the $k$ rows of the matrix, the number of XOR operations performed by the decoder is upper bounded by $\sum_{i=1}^{k}i=\frac{k(k+1)}{2}$. In the back substitution method, the number of XOR operations performed by the decoder at the $i^{th}$ iteration is $k-i$, wherein $i\in[k]$. Therefore, the total number of XOR operations performed at this stage is $\sum_{i=1}^{k}(k-i)=\frac{k(k-1)}{2}$. Adding the number of XOR operations associated with the OFG algorithm as well as the back substitution, we get an upper bound on the computational complexity as $C_{OFG}=k^{2}$. ∎
@@
Next, for the BRH decoder, it is clear that its bootstrap overhead, denoted by $K_{BR}$, is in the range $K_{OFG}\leq K_{BR}\leq K_{BP}$. Furthermore, it is also clear that its decoding failure rate is upper bounded by that of the OFG decoder, as $K_{BR}$ is atleast $K_{OFG}$. When implementing the BRH decoder, choosing values of $K_{BR}$ closer to $K_{OFG}$ leads to achieve a low bootstrap overhead, however, it compromises on computational complexity. The behaviour is vice-versa for choosing $K_{BR}$ closer to $K_{BP}$. Let $\eta_{B}$ be the number of input blocks of an epoch recovered by the BP component of the BRH decoder using $K_{BR}$ droplets. Since $\eta_{B}$ is a discrete random variable that takes values from 00 to $k$, we can numerically evaluate the average value of $\eta_{B}$, denoted by $\operatorname{\mathbb{E}}[\eta_{B}]$ for each value of $K_{BR}$. The remaining $k-\eta_{B}$ blocks will be recovered using the OFG component. Hence, the average computational complexity of the OFG component of the BRH decoder, denoted by $\operatorname{\mathbb{E}}[(k-\eta_{B})^{2}]$ can also be evaluated for each value of $K_{BR}$. With that, the average computational complexity of BRH decoder is given below.
@@
[…]
Computation_hybrid The average computational complexity of the BRH decoder is given by $C_{BR}=K_{BR}\cdot D\cdot\frac{\operatorname{\mathbb{E}}[\eta_{B}]}{k}+\operatorname{\mathbb{E}}[(k-\eta_{B})^{2}]$.
@@
Next, we derive the metrics for the CRH decoder. Recall that the minimum number of input blocks of an epoch that must be recovered exclusively by the BP component of the CRH decoder denoted by $\eta_{c}$, is in the range $0\leq\eta_{c}\leq k$. Note that the bootstrap overhead of the CRH decoder is directly dependent on the $\eta_{c}$ set by the bucket node. When implementing the CRH decoder, it is intuitive that choosing values of $\eta_{c}$ closer to $k$ leads to a lower computational complexity, however it compromises on the bootstrap overhead. In contrast, the behaviour is vice-versa for choosing $\eta_{c}$ closer to 00. Furthermore, it is not straight-forward to determine the exact bootstrap overhead of the CRH decoder, for a given $\eta_{c}$. Nevertheless, for each value of $\eta_{c}\in[k]$, we can numerically evaluate the average bootstrap overhead denoted by $K_{CR}^{av}$, through simulations. As $\eta_{c}$ is fixed, the computational complexity of the OFG component of the CRH decoder will be $(k-\eta_{c})^{2}$, which is a constant. With that, we present the average computational complexity of the CRH decoder as follows.
@@
[…]
Computation_CRH The average computational complexity of the CRH decoder is given by $C_{CR}=K_{CR}^{av}\cdot D\cdot\frac{\eta_{c}}{k}+(k-\eta_{c})^{2}$.
@@
[…]

### v3

@@ 1 Introduction and Background
Blockchain’s distinguished security properties including decentralization, distributed trust, immutability, and transparency are safeguarded by a special class of fully functional nodes, known as full nodes. These nodes serve as the backbone of a blockchain network by replicating and storing all blocks. The key responsibilities of a full node include archiving blockchain history, independently validating new blocks and bootstrapping new nodes joining the network. As a result, full nodes require a large storage space. For instance, the size of the bitcoin blockchain has grown over $554\emph{GB}$ as of March 2024 [9]. Similarly, the XRP ledger size is approximately $26\emph{TB}$ as of July 2023, in which the storage requirement per full node increases by approximately $12\emph{GB}$ per day [10]. As a consequence, scalability of blockchains is limited by its high storage requirements. This causes a significant drop in the number of full nodes, which violates blockchain’s decentralization feature. For instance, the number of operational nodes of Bitcoin blockchain fell from $200,000$ in $2018$ to $47000$ in $2020$ [11]. These numbers clearly indicate that storage is a bottleneck for full nodes in blockchains. Hence, much research effort has been dedicated towards addressing the problem of high storage requirement in blockchains.
@@
Figure 1: Number of contributions that implement RS, Fountain, LDPC and LCC codes to address scalability issues in blockchains (source: [14]).
@@ 1.1 Our Contributions
The security threats proposed in this work inherently capture the vulnerabilities of the vanilla version of the LT coded blockchains [17], wherein decoders based on belief propagation (BP) and on-the-fly Gaussian (OFG) elimination are applicable. Therefore, the impact of our attacks are first presented against such an architecture. In addition, given that we also propose new enhancements to the LT coded framework in the form of BRH and CRH decoders, we also extend the vulnerability study to those enhancements. Overall, our work is the first of its kind in designing optimal decoders for heterogeneous network constraints and also analysing the threats caused by non-oblivious adversaries in LT coded blockchain architecture.
@@ 1.2.1 Blockchain Scalability using Error Correction Codes
Furthermore, as another interesting direction of research, several coded architectures have been proposed to address the scalability issue in blockchains. In this context, authors of [17] present a secure fountain (SeF) architecture for blockchains by using LT codes to provide storage benefits. Ultimately, [17] aims to achieve a near-optimal trade-off between bootstrap overhead and storage savings. Authors of [19] have also worked on similar lines as [17], and suggested raptor codes to be used in place of LT codes to achieve an almost constant decoding complexity, however, with a higher bootstrap overhead. Subsequently, [32] pointed out the problem of high bandwidth requirement for decoding a particular requested block, faced by [17, 19]. To this end, [32] presented two different architectures namely (i) Secure and Private Fountain codes, (ii) Secure and Private Randomly Sampled codes to obtain storage savings, low repair bandwidth per epoch, low communication cost for decoding a requested block, and simultaneously provide privacy of the archival data. However, unlike [17, 19], the architectures in [32] incur higher costs for decoding the complete chain. Another work [33], presents a secure-repair-block protocol based on regenerating codes for sharded blockchains to reduce both storage and bootstrap costs, which makes it easier for a new miner to enter the system. A different study [34] introduces homomorphic hashes in coded blockchains, which allows instantaneous detection of erroneous fragments, thereby avoiding decoding with wrong data. A survey on coded blockchains is also available in [35].
Among the various error correction coding techniques, we adopt LT codes in this work to enhance the scalability, reliability, and data availability of blockchain systems. By fragmenting blockchain data into smaller, redundant coded symbols, LT codes enable efficient data recovery even when some fragments are lost or corrupted, reducing the per-device storage load and enhancing resilience to node failures or network disruptions [36]. Unlike traditional erasure codes, LT codes are non-systematic and do not rely on a fixed generator matrix; each coded block is equally informative, allowing decentralized, load-balanced retrieval from any subset of fragments [37]. This property is particularly beneficial in decentralized environments, as it prevents traffic bottlenecks when multiple nodes simultaneously attempt to sync with the blockchain. Furthermore, LT coding minimizes the need for retransmissions in unreliable networks by enabling decoding from any sufficiently large subset of coded symbols [38], thereby improving transmission efficiency and supporting scalability. As decoding can begin once enough fragments are received, LT-coded systems can also accelerate transaction confirmations, making them well-suited for distributed networks with limited resources.
@@ 1.2.2 Data Availability Attacks in Blockchains
Some preliminary results of this work are available in [48], which captures a part of our contributions on the vulnerabilities of LT coded blockchains to optimized DoS threats. In particular, this work builds upon the work in [48] by providing a comprehensive analysis of LT decoders’ performance metrics and exploring the trade-off between computational complexity and bootstrap overhead, which [48] does not address. While [48] briefly discusses hybrid decoders, this work optimizes them for blockchain decoding at minimal cost under heterogeneous network constraints. In vulnerability analysis, we introduce more advanced and powerful attack strategies and efficient implementation techniques. Specifically, we propose a more effective score-based attack strategy on the BP decoder and devise a computationally efficient algorithm for attacking the OFG decoder, improving upon the attacks presented in [48]. To the best of our knowledge, coded blockchains have not been used in live blockchain deployments, however, they have been validated on real Bitcoin trace data through academic prototypes [36] and studied on IoT networks [51]. While commercial readiness of such architectures are possibly a few years away, we believe our study on the security vulnerabilities is imperative before large-scale deployments as it helps in reducing the risks, prevent downtime, and build more trust and compliance.
@@ 2 LT Coded Blockchain Architecture
To accommodate the rapid growth in the size of a full node, coded blockchain architecture distributes the contents of a full node as coded fragments across multiple smaller-sized storage devices. Fig. 2 shows a visual illustration of coded blockchain architecture. In this architecture, the original blockchain stored on a full node is divided into several epochs, wherein an epoch is defined as a collection of $k$ blocks of the blockchain, for some $k\in\mathbb{Z}^{+}$. For every epoch, a sufficient number of coded blocks referred to as droplets is generated, and these droplets are stored across storage constrained nodes referred to as droplet nodes, as seen in Fig. 2. Henceforth, we assume that a droplet node will store one droplet per epoch, thus storing as many droplets as the number of epochs, as shown clearly in Fig. 2. This way, the storage size of a full node is slashed by distributing it across smaller-sized storage devices with roughly $\frac{1}{k}$ storage capacity. In the next section, we explain how a blockchain is encoded to obtain the droplet nodes using an LT coded architecture [17].
@@ 2.1 Slashing Storage Costs through LT Encoding
When generating the droplets of a full node, LT encoding is done as follows. First, a number $d\in[k]$ is chosen randomly from a suitable probability distribution. Subsequently, $d$ distinct input blocks are uniformly chosen from the $k$ blocks of an epoch, and a bit-wise XOR of the chosen blocks, denoted by $C$, is obtained. Henceforth, this collection of $d$ input blocks of an epoch that are XORed to obtain $C$ is referred to as the neighbours of the droplet $C$. Along with $C$, a binary vector $\mathbf{v}$ of length $k$ is also generated, wherein the $b$-th entry of $\mathbf{v}$ is $1$ if the $b$-th input block is among the $d$ chosen blocks, else the $b$-th entry is 0. This combination of $C$ and $\mathbf{v}$ constitute a droplet for an epoch. An illustration of this procedure is given in Fig. 3, for the case when $k=6$.
@@
Furthermore, we assume that all the droplets of a droplet node that correspond to different epochs, will have the same degree and also the same set of neighbours. In the LT code terminology, the number $d$, which is referred to as the degree, is picked from the popular Robust Soliton Distribution (RSD) [18, Definition 11], denoted by $\Omega_{RS}(\cdot)$ which is defined for a given set of parameters $0<\delta<1$ and $c>0$, as
$\Omega_{RS}(d)=\frac{\rho(d)+\tau(d)}{\beta}\quad\forall d=1,\ldots,k,$ (1)
@@
[…]
@@
$\tau(d)=\begin{cases}\frac{R}{dk}&\text{for }d=1,\ldots,\frac{k}{R}-1\\ \frac{R}{k}\cdot\ln{\left(\frac{R}{\delta}\right)}&\text{for }d=\frac{k}{R}\\ 0&\text{for }d=\frac{k}{R}+1,\ldots,k,\end{cases}$ (3)
@@ 2.2.3 Hybrid Decoder
Similar to the BP and OFG decoders, the bucket node contacts $K\geq k$ droplet nodes and downloads the droplets stored in them. The idea of hybrid decoder, as described in [49, Section III], is to start with the BP decoding process using the $K$ downloaded droplets and retrieve the input blocks of an epoch one by one. Once the BP part of the hybrid decoder exhausts its singletons before retrieving all the $k$ input blocks, the non-singletons remaining in the BP part are used by its OFG counterpart to retrieve the remaining blocks. This idea exploits the advantage of low-complexity decoding of BP decoder by retrieving as many blocks as possible using BP, and also exploits the advantage of the OFG decoder, which guarantees successful decoding by contacting fewer droplet nodes than the BP decoder. Note that the hybrid decoder is said to experience decoding failure if it is unable to recover all $k$ blocks of an epoch from the $K$ droplet nodes despite using its BP and OFG counterparts.
@@ 3.3 Performance Analysis
The average bootstrap overhead of a bucket node employing the BP decoder, to successfully decode an epoch with a probability greater than $1-\delta$ is given by $K_{BP}=k+c\sqrt{k}\ln^{2}{\left(\frac{k}{\delta}\right)}$, where $c>0$ and $0<\delta<1$ are parameters of RSD.
@@
The average computational complexity incurred by the BP decoder for decoding an epoch is given by $C_{BP}=k\cdot\left(\ln\left(\frac{c\cdot k^{\frac{3}{2}}\cdot\ln(k/\delta)}{\delta}\right)+1.577\right)$.
@@
The computational complexity for decoding an epoch can be determined by the number of edges in the bipartite graph $\mathcal{T}$, generated at the start of the BP decoder [17, Appendix B]. This is because, each XOR operation between two droplets corresponds to removal of an edge in the bipartite graph. Subsequently, all the edges of the bipartite graph will be removed at the end of BP decoding process, thereby resulting in as many number of XOR operations. It is important to note that the number of edges in the bipartite graph is random in nature, by the virtue of the random selection of the droplet nodes. Therefore, we calculate its mean value to determine the average computational complexity. In this regard, the average number of edges can be approximated as the product of average bootstrap overhead and the average degree of a droplet. In particular, if $D$ is the average degree of a droplet, according to [18, Theorem 13], we have:
$D=\frac{\sum_{i=1}^{k}i\cdot(\rho(i)+\tau(i))}{\beta}.$ (5)
Recalling the expressions for $\rho(i)$, $\tau(i)$, and $\beta$ as given in Section 2.1, and substituting them into (5), we obtain
$D=\frac{1}{\beta}\left(\frac{1}{k}+\sum_{i=2}^{k}\frac{1}{i-1}+\sum_{i=1}^{\frac{k}{R}-1}\frac{R}{k}+\ln\left(\frac{R}{\delta}\right)\right).$ (6)
Further, (6) can be simplified as follows
$D=\frac{1}{\beta}\left(\sum_{i=2}^{k+1}\frac{1}{i-1}+\sum_{i=1}^{\frac{k}{R}-1}\frac{R}{k}+\ln\left(\frac{R}{\delta}\right)\right).$ (7)
Finally, (7) can be further approximated as follows
$D\approx\frac{1}{\beta}\left(H(k)+1+\ln\left(\frac{R}{\delta}\right)\right),$ (8)
where $H(k)$ is the Harmonic number given by $H(k)=\sum_{n=1}^{k}\frac{1}{n}$. According to [52], for large $k$, the harmonic number can be approximated as $H(k)\approx\ln(k)+\gamma$, where $\gamma\approx 0.577$ is the Euler–Mascheroni constant. Now, substituting $R=c\sqrt{k}\ln{\left(\frac{k}{\delta}\right)}$ and $H(k)=\ln(k)+0.577$ in (8), we obtain the average degree of a droplet as:
$D=\frac{1}{\beta}\cdot\left(\ln\left(\frac{c\cdot k^{\frac{3}{2}}\cdot\ln(k/\delta)}{\delta}\right)+1.577\right).$ (9)
According to [18, Theorem 12], $K_{BP}=k\beta$. With this, we arrive at the average computational complexity as $C_{BP}=K_{BP}D$.
The above two results present closed-form expressions on the average bootstrap overhead and average computational complexity of the BP decoder. Given that the performance of the BP decoder is closely tied with the parameters of the RSD, we observe from (4) and Proposition LABEL:Computation_BP that the bootstrap overhead and the average computational complexity of the BP decoder are inversely proportional to the allowable decoding failure probability $\delta$. Along similar lines, we also observe that both these metrics are directly proportional to the RSD constant $c$.
Now, to derive the bootstrap overhead of OFG decoder, we must investigate the relationship between its probability of decoding failure and the number of downloaded droplets per epoch. To this end, authors of [50] have established an upper bound on the decoding failure probability, denoted by $P_{F}$}, for an LT code whose degree follow any arbitrary degree distribution. As a result, for binary LT codes whose degrees follow RSD, the upper bound in [50, Theorem 1] can be modified as follows:
@@
$\cdot\left[\frac{1}{2}{\color[rgb]{0,0,0}\sum_{d=1}^{k}}\Omega_{RS}(d)\frac{\sum_{l=0}^{d}{w\choose l}{k-w\choose d-l}\left[1-(-1)^{1-l}\right]}{{k\choose d}}\right]^{k(1+\epsilon)},$
where $\epsilon>0$ is the fractional overhead, defined as the ratio of excess droplets (in addition to $k$) and $k$. With that, an upper bound on the bootstrap overhead of the OFG decoder is given in ?THM? LABEL:Overhead_OFG.
@@
Let $\epsilon_{min}$ = $\frac{m}{k}$ where $m\in\mathbb{N}$ such that $P_{F}|_{\epsilon=\epsilon_{min}}\leq\delta$, then the bootstrap overhead of OFG decoder is upper bounded by $K_{OFG}=k(1+\epsilon_{min})$.
@@
Computational complexity of OFG decoder, denoted by $C_{OFG}$ is upper bounded by $k^{2}$.
@@
Computational complexity of the OFG decoder is evaluated as the sum of number of XOR operations performed by the OFG algorithm and back substitution, which together completes the decoding process. In the OFG algorithm, the maximum number of XOR operations associated with the $i^{th}$ droplet, whose binary coefficient entering the partially filled $k\times k$ matrix, is $i$. Hence, to fill the $k$ rows of the matrix, the number of XOR operations performed by the decoder is upper bounded by $\sum_{i=1}^{k}i=\frac{k(k+1)}{2}$. In the back substitution method, the number of XOR operations performed by the decoder at the $i^{th}$ iteration is kik-i, wherein $i\in[k]$. Therefore, the total number of XOR operations performed at this stage is $\sum_{i=1}^{k}(k-i)=\frac{k(k-1)}{2}$. Adding the number of XOR operations associated with the OFG algorithm as well as the back substitution, we get an upper bound on the computational complexity as $C_{OFG}=k^{2}$. ∎
@@
Next, for the BRH decoder, it is clear that its bootstrap overhead, denoted by $K_{BR}$, is in the range $K_{OFG}\leq K_{BR}\leq K_{BP}$. Furthermore, it is also clear that its decoding failure rate is upper bounded by that of the OFG decoder, as $K_{BR}$ is atleast $K_{OFG}$. When implementing the BRH decoder, choosing values of $K_{BR}$ closer to $K_{OFG}$ leads to achieve a low bootstrap overhead, however, it compromises on computational complexity. The behaviour is vice-versa for choosing $K_{BR}$ closer to $K_{BP}$. Let $\eta_{B}$ be the number of input blocks of an epoch recovered by the BP component of the BRH decoder using $K_{BR}$ droplets. Since $\eta_{B}$ is a discrete random variable that takes values from 0 to $k$, we can numerically evaluate the average value of $\eta_{B}$, denoted by $\operatorname{\mathbb{E}}[\eta_{B}]$ for each value of $K_{BR}$. The remaining $k-\eta_{B}$ blocks will be recovered using the OFG component. Hence, the average computational complexity of the OFG component of the BRH decoder, denoted by $\operatorname{\mathbb{E}}[(k-\eta_{B})^{2}]$ can also be evaluated for each value of $K_{BR}$. With that, the average computational complexity of BRH decoder is given below.
@@
The average computational complexity of the BRH decoder is given by $C_{BR}=K_{BR}\cdot D\cdot\frac{\operatorname{\mathbb{E}}[\eta_{B}]}{k}+\operatorname{\mathbb{E}}[(k-\eta_{B})^{2}]$, in which $K_{OFG}\leq K_{BR}\leq K_{BP}$ and $\eta_{B}\in\{0,1,2,\dots,k\},\ \text{where }k\in\mathbb{Z}^{+}$.
@@
Next, we derive the metrics for the CRH decoder. Recall that the minimum number of input blocks of an epoch that must be recovered exclusively by the BP component of the CRH decoder denoted by $\eta_{c}$, is in the range $0\leq\eta_{c}\leq k$. Note that the bootstrap overhead of the CRH decoder is directly dependent on the $\eta_{c}$ set by the bucket node. When implementing the CRH decoder, it is intuitive that choosing values of $\eta_{c}$ closer to $k$ leads to a lower computational complexity, however it compromises on the bootstrap overhead. In contrast, the behaviour is vice-versa for choosing $\eta_{c}$ closer to 0. Furthermore, it is not straight-forward to determine the exact bootstrap overhead of the CRH decoder, for a given $\eta_{c}$. Nevertheless, for each value of $\eta_{c}\in[k]$, we can numerically evaluate the average bootstrap overhead denoted by $K_{CR}^{av}$, through simulations. As $\eta_{c}$ is fixed, the computational complexity of the OFG component of the CRH decoder will be $(k-\eta_{c})^{2}$, which is a constant. With that, we present the average computational complexity of the CRH decoder as follows.
@@
The average computational complexity of the CRH decoder is given by $C_{CR}=K_{CR}^{av}\cdot D\cdot\frac{\eta_{c}}{k}+(k-\eta_{c})^{2}$, where $0\leq\eta_{c}\leq k$.
@@
Explanation follows from the proof of ?THM? LABEL:Computation_hybrid. ∎